Suppose you bought a bond with an annual coupon rate of 8 percent one year ago for $880. The bond sells for $910 today. 

a. Assuming a $1,000 face value, what was your total dollar return on this investment over the past year? 

Total dollar return = $_______ 

b. What was your total nominal rate of return on this investment over the past year? 

 Nominal rate of return = _______ % 

c. If the inflation rate last year was 3 percent, what was your total real rate of return on this investment? 

 Real rate of return = _______ %
Answer #1
a] Total dollar return = Total interest + Increase in price = 80+(910-880) = $                 110
b] Total nominal rate of return = 110/880 = 12.50%
c] Total real rate of return = 1.125/1.03-1 = 9.22%